OpenStudy (anonymous):

convert polar to rectangular: r=8cos(theta)-10sin(theta) r^2=8rcos(theta)-10rsin(theta) then does it go into x^2+8x and y^2-10y and complete the square? no idea if thats right...

OpenStudy (anonymous):

yea just leave like that imo o-o..

OpenStudy (anonymous):

so (x+4)^2+(y-5)^2=41?

OpenStudy (anonymous):

r=radical 41

OpenStudy (anonymous):

wait a bit

OpenStudy (anonymous):

x^2 +y^2 = 8x - 10y x^2 - 8x = -y^2 - 10y -25 + 25 (x-4^)2 -16 =- (y+5)^2 + 25 (x-4)^2 + (y-5)^2 = 41 k now i'm 100% positive that this's right

OpenStudy (anonymous):

shouldnt it be x+4? because its x^2+8 and 8/2=4, 4^2=16

OpenStudy (anonymous):

just kidding i got it.

OpenStudy (anonymous):

its y+5